For the sequence a sub n = 7/3^n, find its fifth partial sum and its nth partial sum.
Help would truly be appreciated!

Answer by rothauserc(4718)

I need to know if you are asking about
a(sub n) = 7 / (3^n) or
a(sub n) = (7/3)^n
ok
you answered a(sub n) = 7 / 3^n
note that
a(sub n) = 7 * ( 1 / 3^n)
( 1 / 3^n) is a geometric series with a = 1/3 and r = 1/3
therefore
sum of n terms = 7 * a * ((1 - r^n) / (1 - r))
sum of first 5 terms = 7 * (1/3) * ((1 - (1/3)^5) / ( 1 - (1/3))
sum of first 5 terms = 3.485596708